OpenStudy (kayders1997):

Will someone help me with this find the integral of (x^2+1)/square root of x

OpenStudy (steve816):

\[\large \int\limits_{}^{}\frac{ x^2+1 }{ \sqrt{x} }~dx=\int\limits_{}^{}\frac{ x^2 }{ x^\frac{ 1 }{ 2 } }+\frac{ 1 }{ x^\frac{ 1 }{ 2 }}~dx\]

OpenStudy (steve816):

Ok now, you can simplify the fractions ya?

OpenStudy (kayders1997):

ohhh right, I forgot that you can split fractions, so it would be x^3/2+1/square root of x

OpenStudy (kayders1997):

@harman.singh

OpenStudy (harman.singh):

Is that your answer for the whole integral or one fraction?

OpenStudy (kayders1997):

whole

OpenStudy (kayders1997):

@harman.singh

OpenStudy (kayders1997):

@zepdrix

zepdrix (zepdrix):

whut :U

OpenStudy (kayders1997):

what?

OpenStudy (kayders1997):

I may just be overthinking

zepdrix (zepdrix):

`so it would be x^3/2+1/square root of x` \[\large\rm \int\limits x^{3/2}+x^{-1/2}~dx\]Yes, before integrating.

OpenStudy (kayders1997):

ok

zepdrix (zepdrix):

You figure out the rest of it lady Kay? :d

OpenStudy (kayders1997):

yes 2/5x^5/2+2x^1/2

zepdrix (zepdrix):

yay good job

OpenStudy (kayders1997):

\[2/5x^2\sqrt{x}+2\sqrt{x}+c\]

OpenStudy (kayders1997):

:)

OpenStudy (kayders1997):

I forgot you could split the fractions I was trying to do u sub lol

zepdrix (zepdrix):

Bahhh :d

OpenStudy (kayders1997):

lol what?